# Marginal treatment effects in the absence of instrumental variables

### v1

@@ A.2 Further discussion on the nonlinearity assumption
When $\dim\left(X^{C}\right)=1$, taking the derivative of both sides of (A.1) and evaluating at $x^{C}$, we obtain
$m_{d0}^{\left(1\right)}\left(x^{C}\right)=\beta_{d}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)\pi_{0}^{\left(1\right)}\left(x^{C}\right)=\beta_{d}^{C},$
which identifies $\beta_{d}^{C}$ for $d=0,1$. The remaining part of the proof is the same as the proof of Theorem 1.
Compared with Assumption NL1, the condition of Theorem A.1 may be satisfied even if $\pi_{0}\left(x^{C}\right)$ is a strictly monotonic function as long as it has a stationary point, such as $\pi_{0}\left(x^{C}\right)=\left(x^{C}\right)^{3}$ at $x^{C}=0$. The two theorems below show that the local irrelevance condition can be extended to the case of multiple continuous covariates.
Theorem A.2.
@@
$\partial_{k}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,k}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{k}\pi_{0}\left(\breve{x}^{C}\right),$ (A.4)
$\partial_{j}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,j}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{j}\pi_{0}\left(\breve{x}^{C}\right).$ (A.5)
Since $\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)\neq 0$ by condition (ii), we can find $g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)$ from (A.4) to be
@@
and substitute it into (A.5) to identify $\beta_{d,j}^{C}$ as
@@
The identification of $\beta_{d,k}^{C}$ is analogous to Theorem A.1. Then, the identification of all the other coefficients $\beta_{d,j}^{C}$ in $\beta_{d}^{C}$ and the remaining components of the model is the same as in Theorem A.2 and Theorem 1, respectively.
Alternatively, the identification of coefficients of continuous covariates can also be attained by exploiting the local irrelevance of the function $g_{d}\left(p\right)$.
Theorem A.4.
Theorem 1 holds if Assumption NL is replaced by that, there exists $x^{C}$ in the support of $X^{C}$ such that $g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)=0$ for $d=0,1$, and functions $\pi_{0}$ and $m_{d0}$ are differentiable at $x^{C}$.
@@
The above discussion further illustrates the overidentification feature of our strategy. In fact, the identifying power of our method is determined by the number of nonlinearity points in the data satisfying Assumption NL or Theorems A.1-A.4.
A.3 Limited valued outcomes
@@
Assumption A.L (Linear index). Assume that $Y_{d}=l\left(X^{\prime}\gamma_{d},V_{d}\right)$, $d=0,1$, holds for some known link function $l$ and unknown coefficients $\gamma_{d}$, where $V_{d}$ is unobserved error term of $Y_{d}$.
The generalized linear model applies to a variety of frequently encountered limited dependent variables, such as $l\left(t,v\right)=1\left\{t\geq v\right\}$ for binary $Y\in\left\{0,1\right\}$, $l\left(t,v\right)=\max\left(t-v,0\right)$ for censored (at zero) $Y\geq 0$, $l\left(t,v\right)=\exp\left(t+v\right)\left/\left[1+\exp\left(t+v\right)\right]\right.$ for proportion-valued $Y\in\left(0,1\right)$, $l\left(t,v\right)=\sum_{j=1}^{s}1\left\{t+v\geq c_{j}\right\}$ for ordered $Y\in\left\{0,1,\cdots,s\right\}$ with known thresholds $c_{1}<c_{2}<\cdots<c_{s}$, and so forth. Although setting $l\left(t,v\right)=t+v$ reduces to the linear additive model, the following discussion does not include Theorem 1 as a special case because $\gamma_{d}$ here can be only identified up to scale for a general link function $l$. Any changes in the scaling of $\gamma_{d}$ can be freely absorbed into $l$, and a scale normalization is needed for identification of $\gamma_{d}$ and $l$.
Assumption A.N (Normalization). Decompose $X^{\prime}\gamma_{d}=X^{C\prime}\gamma_{d}^{C}+X^{D\prime}\gamma_{d}^{D}$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. Assume that $\gamma_{d,1}^{C}$, the first element of $\gamma_{d}^{C}$, equal to 1.
Assumption A.N imposes the convenient normalization that the first continuous covariate has a unit coefficient. This scaling of $\gamma_{d}$ is arbitrary and is innocuous because our focus is on the identification of MTE, which is a difference in the conditional expectations of $l\left(X^{\prime}\gamma_{d},V_{d}\right)$, rather than on separate identification of $\gamma_{d}$ and $l$. Since in the generalized linear model the potential outcomes are not necessarily additive in the error term, the mean independence assumption needs to be strengthened to a stricter distributional independence assumption.
Assumption A.CI (Conditional Independence). Assume that $\left.V_{d}\perp\!\!\!\!\perp X\right|V$ for $d=0,1$, namely, $V_{d}$ is independent of $X$ conditional on $V$, where $V$ is the reduced-form treatment error in equation (3).
Recalling that $V\perp\!\!\!\!\perp X$ by definition, Assumption A.CI is equivalent to the full independence $\left(V_{d},V\right)\perp\!\!\!\!\perp X$ for $d=0,1$, which implies that both the marginal distribution of $V_{d}$ and the copula of $V_{d}$ and $V$ are independent of $X$. Under Assumptions A.L and A.CI, we have
@@
The additive nonseparability of the observables and unobservables constitutes the primary difficulty in identifying MTE in the limited outcome case. Meanwhile, Assumptions A.L and A.CI lead to a double index form of the observable outcome regression functions for each treatment status, that is,
$E\left[\left.Y1\left\{D=d\right\}\right|X=x\right]=r_{d}\left(x^{\prime}\gamma_{d},\pi\left(x\right)\right)$ (A.6)
@@
$\Delta^{\text{MTE}}\left(x,v\right)=\partial_{2}r_{1}\left(x^{\prime}\gamma_{1},v\right)+\partial_{2}r_{0}\left(x^{\prime}\gamma_{0},v\right),$ (A.9)
@@
Assumption A.NL (Non-Linearity). Assume that the functions $\pi_{0}$ and $r_{d}$ are differentiable and denote their partial derivatives with respect to the $k$-th argument as $\partial_{k}\pi_{0}$ and $\partial_{k}r_{d}$, where $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$ and $r_{d}$, $d=0,1$, are defined in (A.6) and (A.6). Assume that there exist two vectors $x^{C}$ and $\tilde{x}^{C}$ on the support of $X^{C}$ and two elements $k$ and $j$ of the set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that (i) $\partial_{1}r_{d}\left(x^{C\prime}\gamma_{d}^{C},\pi_{0}\left(x^{C}\right)\right)\neq 0$, (ii) $\partial_{1}r_{d}\left(\tilde{x}^{C\prime}\gamma_{d}^{C},\pi_{0}\left(\tilde{x}^{C}\right)\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,k}^{C}$, (iv) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,j}^{C}$, (v) $\partial_{l}\pi_{0}\left(\tilde{x}^{C}\right)\neq\partial_{1}\pi_{0}\left(\tilde{x}^{C}\right)\gamma_{d,l}^{C}$ for $l=2,\cdots,\dim\left(X^{C}\right)$, and
@@
Assumption A.NL (i)-(ii) require $r_{d}$ to depend on the linear index, and (iii)-(vi) essentially require some nonlinear variation in $\pi_{0}$ under the scale normalization. As in Assumption NL, it is difficult to construct examples other than $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ that violates Assumption A.NL. Finally, a support assumption and an invertibility assumption are imposed for technical reasons.
Assumption A.S (Support). For each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, assume for some $x_{k}^{D}\neq 0$ in the support of $X_{k}^{D}$ that there exists $x^{C}\left(k\right)$ in the support of $X^{C}$ such that $\pi\left(x^{C}\left(k\right),x^{Dk}\right)$ is in the support of $\pi_{0}\left(X^{C}\right)$ and that $x^{C}\left(k\right)^{\prime}\gamma_{d}^{C}+x_{k}^{D}\gamma_{d,k}^{D}$ is in the support of $X^{C\prime}\gamma_{d}^{C}$ for $d=0,1$.
Assumption A.I (Invertibility). Assume that the function $r_{d}$ is invertible on its first argument for $d=0,1$.
@@
Theorem A.5.
If Assumptions A.L, A.NL, A.CI, A.S, A.N, and A.I hold, then $\gamma_{d}$ and $r_{d}\left(t,p\right)$ at all points in the support of $X^{\prime}\gamma_{d}$ and $P$ are identified for $d=0,1$.
Having established identification of $\gamma_{d}$ and $r_{d}$, the MTE for limited valued outcomes can be identified by (A.9) and estimated by the semiparametric least squares method of Escanciano et al. (2016). Alternatively, the LIV estimation procedure may be adapted according to
$E\left[\left.Y\right|X=x\right]=s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},\pi\left(x\right)\right),$
where
$s\left(t_{0},t_{1},p\right)=r_{0}\left(t_{0},p\right)+r_{1}\left(t_{1},p\right),$
@@
$\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $\frac{s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},v_{2}\right)-s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},v_{1}\right)}{v_{2}-v_{1}}.$

### v2

@@ Appendix B Further discussion on the nonlinearity assumption
When $\dim\left(X^{C}\right)=1$, it follows from (A.1) and the assumption of Theorem B.1 that
Hence, $\beta_{d}^{C}$ is identified as
The remaining part of the proof is the same as the proof of Theorem 1.
The assumption of Theorem B.1 in place of NL1 requires the univariate function $\pi_{0}$ to be not one-to-one, but imposes no other smoothness or continuity restriction on $\pi_{0}$. Therefore, it doesn’t require $X^{C}$ to be continuously valued. In the extreme case of $X^{C}$ containing only one binary covariate, this assumption will be equivalent to the full irrelevance of $X^{C}$ to the treatment probability, which is a condition suggested by Chamberlain (1986) for identification of the sample selection model. The following two theorems show that Assumption NL1 and the assumption of Theorem B.1 can be extended to the case of multiple continuous covariates.
Theorem B.2.
@@
$\partial_{k}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,k}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{k}\pi_{0}\left(\breve{x}^{C}\right),$ (B.1)
$\partial_{j}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,j}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{j}\pi_{0}\left(\breve{x}^{C}\right).$ (B.2)
Since $\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)\neq 0$ by condition (ii), we can find $g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)$ from (B.1) to be
@@
and substitute it into (B.2) to identify $\beta_{d,j}^{C}$ as
@@
The identification of $\beta_{d,k}^{C}$ is analogous to Theorem 1. Then, the identification of all the other coefficients $\beta_{d,j}^{C}$ in $\beta_{d}^{C}$ and the remaining components of the model is the same as in Theorem B.2 and Theorem 1, respectively.
Alternatively, the identification of coefficients of continuous covariates can be attained by exploiting the local irrelevance of the function $g_{d}\left(p\right)$.
Theorem B.4.
Theorem 1 holds if Assumption NL is replaced by the existence of a vector $x^{C}$ in the support of $X^{C}$ such that $g_{d}^{\left(1\right)}\left(\pi_{0}\left(x^{C}\right)\right)=0$ for $d=0,1$, and that functions $\pi_{0}$ and $m_{d0}$ are differentiable at $x^{C}$.
@@
The above discussion further illustrates the overidentification feature of our strategy. In fact, the identifying power of our method is determined by the number of nonlinearity points in the data satisfying Assumption NL or Theorems B.1-B.4, given different benchmark values of $X^{D}$.
Appendix C Limited valued outcomes
@@
Assumption C.L (Linear index). Assume that $Y_{d}=l\left(X^{\prime}\gamma_{d},V_{d}\right)$, $d=0,1$, holds for some known link function $l$ and unknown coefficients $\gamma_{d}$, where $V_{d}$ is unobserved error term of $Y_{d}$.
The generalized linear model applies to a variety of frequently encountered limited dependent variables, such as $l\left(t,v\right)=1\left\{t\geq v\right\}$ for binary $Y\in\left\{0,1\right\}$, $l\left(t,v\right)=\max\left(t-v,0\right)$ for censored (at zero) $Y\geq 0$, $l\left(t,v\right)=\exp\left(t+v\right)\left/\left[1+\exp\left(t+v\right)\right]\right.$ for proportion-valued $Y\in\left(0,1\right)$, $l\left(t,v\right)=\sum_{j=1}^{s}1\left\{t+v\geq c_{j}\right\}$ for ordered $Y\in\left\{0,1,\cdots,s\right\}$ with known thresholds $c_{1}<c_{2}<\cdots<c_{s}$, and so forth. Although setting $l\left(t,v\right)=t+v$ reduces to the linear additive model, the following discussion doesn’t include Theorem 1 as a special case because $\gamma_{d}$ here can be only identified up to scale for a general link function $l$. Any changes in the scaling of $\gamma_{d}$ can be freely absorbed into $l$, and a scale normalization is needed for identification of $\gamma_{d}$ and $l$.
Assumption C.N (Normalization). Decompose $X^{\prime}\gamma_{d}=X^{C\prime}\gamma_{d}^{C}+X^{D\prime}\gamma_{d}^{D}$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. Assume that $\gamma_{d,1}^{C}$, the first element of $\gamma_{d}^{C}$, equal to 1.
Assumption C.N imposes the convenient normalization that the first continuous covariate has a unit coefficient. This scaling of $\gamma_{d}$ is arbitrary and is innocuous because our focus is on the identification of MTE, which is a difference in the conditional expectations of $l\left(X^{\prime}\gamma_{d},V_{d}\right)$, rather than on separate identification of $\gamma_{d}$ and $l$. Since in the generalized linear model the potential outcomes are not necessarily additive in the error term, the mean independence assumption needs to be strengthened to a stricter distributional independence assumption.
Assumption C.CI (Conditional Independence). Assume that $\left.V_{d}\perp\!\!\!\!\perp X\right|V$ for $d=0,1$, namely, $V_{d}$ is independent of $X$ conditional on $V$, where $V$ is the reduced-form treatment error in equation (3).
Recalling that $V\perp\!\!\!\!\perp X$ by definition, Assumption C.CI is equivalent to the full independence $\left(V_{d},V\right)\perp\!\!\!\!\perp X$ for $d=0,1$, which implies that both the marginal distribution of $V_{d}$ and the copula of $V_{d}$ and $V$ are independent of $X$. Under Assumptions C.L and C.CI, we have
@@
The additive nonseparability of the observables and unobservables constitutes the primary difficulty in identifying MTE in the limited outcome case. Meanwhile, Assumptions C.L and C.CI lead to a double index form of the observable outcome regression functions for each treatment status, that is,
$E\left[\left.Y1\left\{D=d\right\}\right|X=x\right]=r_{d}\left(x^{\prime}\gamma_{d},\pi\left(x\right)\right)$ (C.1)
@@
$\Delta^{\text{MTE}}\left(x,v\right)=\partial_{2}r_{1}\left(x^{\prime}\gamma_{1},v\right)+\partial_{2}r_{0}\left(x^{\prime}\gamma_{0},v\right),$ (C.4)
@@
Assumption C.NL (Non-Linearity). Assume that the functions $\pi_{0}$ and $r_{d}$ are differentiable and denote their partial derivatives with respect to the $k$-th argument as $\partial_{k}\pi_{0}$ and $\partial_{k}r_{d}$, where $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$ and $r_{d}$, $d=0,1$, are defined in (C.1) and (C.1). Assume that there exist two vectors $x^{C}$ and $\tilde{x}^{C}$ on the support of $X^{C}$ and two elements $k$ and $j$ of the set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that (i) $\partial_{1}r_{d}\left(x^{C\prime}\gamma_{d}^{C},\pi_{0}\left(x^{C}\right)\right)\neq 0$, (ii) $\partial_{1}r_{d}\left(\tilde{x}^{C\prime}\gamma_{d}^{C},\pi_{0}\left(\tilde{x}^{C}\right)\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,k}^{C}$, (iv) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,j}^{C}$, (v) $\partial_{l}\pi_{0}\left(\tilde{x}^{C}\right)\neq\partial_{1}\pi_{0}\left(\tilde{x}^{C}\right)\gamma_{d,l}^{C}$ for $l=2,\cdots,\dim\left(X^{C}\right)$, and
@@
Assumption C.NL.(i)-(ii) require $r_{d}$ to depend on the linear index, and (iii)-(vi) essentially require some nonlinear variation in $\pi_{0}$ under the scale normalization. As in Assumption NL, it is difficult to construct examples other than $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ that violates Assumption C.NL. Finally, a support assumption and an invertibility assumption are imposed for technical reasons.
Assumption C.S (Support). For each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, assume for some $x_{k}^{D}\neq 0$ in the support of $X_{k}^{D}$ that there exists $x^{C}\left(k\right)$ in the support of $X^{C}$ such that $\pi\left(x^{C}\left(k\right),x^{Dk}\right)$ is in the support of $\pi_{0}\left(X^{C}\right)$ and that $x^{C}\left(k\right)^{\prime}\gamma_{d}^{C}+x_{k}^{D}\gamma_{d,k}^{D}$ is in the support of $X^{C\prime}\gamma_{d}^{C}$ for $d=0,1$.
Assumption C.I (Invertibility). Assume that the function $r_{d}$ is invertible on its first argument for $d=0,1$.
@@
Theorem C.1.
If Assumptions C.L, C.NL, C.CI, C.S, C.N, and C.I hold, then $\gamma_{d}$ and $r_{d}\left(t,p\right)$ at all points in the support of $X^{\prime}\gamma_{d}$ and $P$ are identified for $d=0,1$.
Having established identification of $\gamma_{d}$ and $r_{d}$, the MTE for limited valued outcomes can be identified by (C.4) and estimated by the semiparametric least squares method of Escanciano et al. (2016). Alternatively, the LIV estimation procedure may be adapted according to
$E\left[\left.Y\right|X=x\right]=s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},\pi\left(x\right)\right)$
@@
$\Delta^{\text{LATE}}\left(x,v_{1},v_{2}\right)$ $=$ $\frac{s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},v_{2}\right)-s\left(x^{\prime}\gamma_{0},x^{\prime}\gamma_{1},v_{1}\right)}{v_{2}-v_{1}},$
where
$s\left(t_{0},t_{1},p\right)=r_{0}\left(t_{0},p\right)+r_{1}\left(t_{1},p\right).$